Julie inherited $10000 from her uncle . She invested part in stocks and the rest in bonds. The stocks paid an annual interest rate of 3% and the bonds paid 5% annually. If her total annual income from these two investment was $413. How much did she invest in each?
Answer by checkley77(12569)

.05x+.03(10,000-x)=413
.05x+300-.03x=413
.02x=413-300
.02x=113
x=113/.02
x=5,650 amount invested @ 5%.
10,000-5,650=4,350 amount invested @ 3%.
Proof:
.05*5,650+.03*4,350=413
282.5+130.5=413
413=413